OpenStudy (anonymous):

an investor invested a total of 2,900 in two mutual funds and one fund earned a 9% profit while the other earned a 7% profit,if the investor total profit was 233. how much was invested in each mutual profit? I took 9 divides by 2900 = 322.2 is this right? then i divided 7 by 233= 33.3 is this right

OpenStudy (anonymous):

i will appreciate anyone that can help me?

OpenStudy (anonymous):

To hard

jimthompson5910 (jim_thompson5910):

let x and y be the two amounts

jimthompson5910 (jim_thompson5910):

x be the amount the gets you a 9% profit and y be the amount that gets you the 7% profit

jimthompson5910 (jim_thompson5910):

so this means x+y = 2900 is your first equation

jimthompson5910 (jim_thompson5910):

solve for y to get y = 2900 - x

jimthompson5910 (jim_thompson5910):

since x dollars invested gives you a 9% profit, this means we go from x dollars to 1.09x dollars y dollars invested gives you a 7% profit, this means we go from y dollars to 1.07y dollars they add up to 1.09x + 1.07y

jimthompson5910 (jim_thompson5910):

it's given that the "total profit was 233", so 1.09x + 1.07y = 233 Now plug in y = 2900 - x to get 1.09x + 1.07y = 233 1.09x + 1.07(2900 - x) = 233 and solve for x

jimthompson5910 (jim_thompson5910):

oh wait...hmm i'm thinking of something else, the second equation should be 0.09x + 0.07y = 233 plug in y = 2900 - x to get 0.09x + 0.07(2900 - x) = 233

jimthompson5910 (jim_thompson5910):

now solve 0.09x + 0.07(2900 - x) = 233 for x

OpenStudy (anonymous):

u add 0.09 + 0.07 = 0.16 then u x 2900 =464 so 233x = 464?

OpenStudy (anonymous):

i have to get the profit for 9% and also 7 % so I dont think I got it right?

jimthompson5910 (jim_thompson5910):

0.09x + 0.07(2900 - x) = 233 0.09x + 0.07(2900) - 0.07x = 233 0.09x + 203 - 0.07x = 233 0.02x + 203 = 233 0.02x = 233 - 203 0.02x = 30 x = 30/0.02 x = 1500 So $1500 was invested and it generated a 9% profit

jimthompson5910 (jim_thompson5910):

use this x value to find y

OpenStudy (anonymous):

thank you I got it. thank you so much you helped me a lot. you should honestly be a tutor, to bad you dont live in Boise idaho I would pay you to tutor me.

jimthompson5910 (jim_thompson5910):

I am a tutor lol, I could tutor online if you want

jimthompson5910 (jim_thompson5910):

that's kinda what I just did right now